OpenStudy (anonymous):

Find the value of the circumference of a circle with the given diameter. 3.21 ft. C= a. 1.6 pi b. 3.21 pi c. 6.42 pi

jimthompson5910 (jim_thompson5910):

You use this formula Circumference = pi*diameter C = pi*d

OpenStudy (anonymous):

C= pi * 3.21 C= 3.14 * 3.21 C= 10.07

OpenStudy (anonymous):

Am I right?

OpenStudy (anonymous):

circumference is not measured in radians. why are the possible answers measured in radians. @jim_thompson5910 right

OpenStudy (ryan123345):

Circumference = diameter x 3.14

jimthompson5910 (jim_thompson5910):

leave it in terms of pi, so it's just 3.21pi

OpenStudy (anonymous):

I did that Ryan

jimthompson5910 (jim_thompson5910):

the answer you got is an approximate answer, 3.21pi is an exact answer

OpenStudy (anonymous):

are a, b, and c possible measure for pi or possible circumferences

jimthompson5910 (jim_thompson5910):

possible circumferences

OpenStudy (anonymous):

wouldn't that be measured in ft?

OpenStudy (anonymous):

Yes

jimthompson5910 (jim_thompson5910):

yeah it doesn't say, but I'm assuming they're implying 1.6 pi ft, 3.21 pi ft, etc...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I'm sorry it's not feet it doesn't say way it is all it says are the numbers.

OpenStudy (anonymous):

What*

OpenStudy (ryan123345):

3.21pi

OpenStudy (anonymous):

Are you sure?

OpenStudy (ryan123345):

lol.. trust me

OpenStudy (ryan123345):

thats 7th grade stuff

jimthompson5910 (jim_thompson5910):

yes it's 3.21pi

OpenStudy (anonymous):

How do I choose best answer when I'm on a iphone ? LOLOLOL

OpenStudy (ryan123345):

then you call me and ask me :p

OpenStudy (anonymous):

Well, they're teaching it in 10th now lol

OpenStudy (ryan123345):

yes they are

OpenStudy (anonymous):

Lol thank you (: ill choose best answer when I get on my computer (:

OpenStudy (ryan123345):

haha, ok thank you so much :D